Proof whether the following operations are inner product operations:

⟨x, y⟩ = 2x1y1 − x1y2 − x2y1 + 2x2y2, x=(x1, x2), y=(y1, y2)



1
Expert's answer
2021-11-17T07:10:01-0500

We must verify linearity,symmetry and positive definiteness properties.

We can define <x,y> through matrix operations as

"<x,y>=x^T\cdot A\cdot y, \space x=\begin{pmatrix}

 x_1 \\

 x_2\end{pmatrix},x^T=(x_1\space x_2),y=\begin{pmatrix}

 y_1 \\

 y_2\end{pmatrix},

A=\begin{pmatrix}

 2 & -1 \\

 -1 & 2

\end{pmatrix}"


and use linearity of transposition operation "^T" and matrix multiplication

So we have

"<x^{(1)}+x^{(2)},y>=(x^{(1)}+x^{(2)})^T\cdot A\cdot y=\\
( (x^{(1)})^T+(x^{(2)}))\cdot A\cdot y=((x^{(1)})^T\cdot A+(x^{(2)})^T\cdot A\cdot y=\\
(x^{(1)})^T\cdot A\cdot y+(x^{(2)})^T\cdot A\cdot y=<x^{(1)},y>+<x^{(2)},y>"

and

"<c\cdot x,y>=(c\cdot x)^T\cdot A \cdot y=(c\cdot x^T)\cdot A\cdot y=\\
c\cdot (x^T\cdot A\cdot y)=c\cdot <x,y>"

Thus linearity by the first argument of <,> is proved.

2) Symmetry property

"<x,y>=x^T\cdot A \cdot y=[x^T\cdot A \cdot y\in R]=(x^T\cdot A \cdot y)^T=\\
y^T\cdot A^T\cdot (x^T)^T=[A^T=A,\space A-symmetric,(x^T)^T=x]=y^T\cdot A\cdot x=<y,x>"

From symmetry of <x,y> linearity by 2-argument is proved also.

3) Positive definiteness property.

This property better to verify by using simplest definition of <x,y>.

Let "x\ne\overline 0" or "x_1^2+x_2^2> 0" . We have

<x,x>="2x_1^2-2x_1x_2+2x_2^2=(x_1-x_2)^2+x_1^2+x_2^2>0"

Thus all properties of inner product is proved.